a reaction of second order with respect to a reactant.how is the rate of reaction affected if the concentration of the reactants is reduced half? what is the unit of rate constant of such reaction?
1
Expert's answer
2012-09-14T08:54:05-0400
Look in your text book for the 2nd order chemical reaction (shown below) and plug n chug fake numbers into [A], the concentration, t (time in seconds), and
A0 and see what happens to k (the rate of the reaction) when you double and reduce A by half.

1/[A] = kt + 1/A0

Rate = k[A]^2

mol/sec = k(mol/L)^2 = k(mol2/L2)

k = L2/mol-sec